OpenStudy (anonymous):

Select the inequality that corresponds to the given graph. graph of an inequality with a dashed line through the points negative 2 comma 0 and 0 comma 4 and shading above the line 2x - 5y -10 x - y > 3 2x - y < -4 x + 2y -8

OpenStudy (mathstudent55):

What are choices 1 and 4? You left out the inequality sign.

OpenStudy (anonymous):

\[2x - 5y \ge -10 \]

OpenStudy (anonymous):

\[x + 2y \ge -8 \]

OpenStudy (mathstudent55):

Ok, thanks. The dashed line goes through (-2, 0) and (0, 4) Its slope is: (0 - 4)/(2 - 0) = -4/-2 = 2 Its equation is: y - 0 = 2(x + 2) y = 2x + 4 2x - y = -4 Answer can only be 2x - y < -4 or 2x - y > -4